# DELVE & MAGIC Collaborations Spectroscopic Analysis of Pictor II: a very low metallicity ultra-faint dwarf galaxy bound to the Large Magellanic Cloud⋆

### v1

@@ Abstract
The advent of wide and deep optical surveys has precipitated the discovery of a large number of ultra-faint dwarf galaxies (UFDs, ) around the Milky Way (MW) (e.g., Willman et al., 2005; Belokurov et al., 2007; Laevens et al., 2015; Bechtol et al., 2015; Koposov et al., 2015a; Drlica-Wagner et al., 2015; Cerny et al., 2023b; Smith et al., 2023; Homma et al., 2024; Tan et al., 2025). UFDs are among the oldest, most chemically pristine, and dark matter dominated galaxies known (McConnachie, 2012; Simon, 2019; Pace, 2024) and are excellent probes to test CDM at small scales (Bullock & Boylan-Kolchin, 2017). UFDs are old and metal-poor, and their stars trace chemical enrichment from the first Population III stars (Simon & Geha, 2007; Brown et al., 2014; Frebel & Norris, 2015). The population of UFDs probes galaxy formation before the epoch of reionization (e.g., Bose et al., 2018) and constrains the small-scale power spectrum of dark matter (e.g., Jethwa et al., 2018; Newton et al., 2018; Nadler et al., 2021).
@@
The Magellanic SatelLites Survey111MagLiteS was a precursor survey to the DECam Local Volume Exploration survey. (MagLiteS; PI: K. Bechtol) was designed to search for additional satellites near the Magellanic Clouds (Drlica-Wagner et al., 2016), particularly on the southern side of the LMC/SMC that the DES did not cover. Three UFD candidates were discovered with MagLiteS: Pictor II (Pic II), Carina II (Car II), and Carina III (Car III) (Drlica-Wagner et al., 2016; Torrealba et al., 2016a). The radial velocities of Car II and Car III (Li et al., 2018a) suggested they were associated with the LMC based on simulated LMC satellites (Jethwa et al., 2016) and this association was confirmed with full 6-D phase space information from Gaia DR2 proper motion measurements (Gaia Collaboration et al., 2016, 2018) and orbit models that included the LMC (Erkal & Belokurov, 2020; Patel et al., 2020). Several other UFDs, Horologium I (Hor I), Hydrus I (Hyi I), Phoenix II (Phe II), and Reticulum II (Ret II), have additionally been associated with the LMC (Koposov et al., 2018; Simon, 2018; Kallivayalil et al., 2018; Erkal & Belokurov, 2020; Patel et al., 2020; Correa Magnus & Vasiliev, 2022; Battaglia et al., 2022; Pace et al., 2022; Vasiliev, 2023).
The LMC UFD population enables studying UFDs that formed in a different and lower mass environment compared to the in-situ MW UFDs. Star formation history measurements suggest that the LMC satellites were quenched more recently compared to MW satellites (Sacchi et al., 2021; Durbin et al., 2025). This may be due to patchy reionization allowing lower density environments to form stars longer (e.g., Lunnan et al., 2012; Katz et al., 2020; Kim et al., 2023). Conversely, no difference has been seen in the size-luminosity plane between the LMC and MW UFDs (Richstein et al., 2024) but the number of LMC confirmed UFDs is an order of magnitude smaller than MW UFDs and confirming more LMC UFDs is key to studying environmental differences.
@@ 2.1.2 Velocity and Metallicity Measurements
For our primary analysis, we combined repeat measurements at the catalog level. For the combined velocities we fit a free velocity and dispersion to the repeat measurements and use the mean velocity and error for the star’s intrinsic properties. This method has the appealing limits that it will match the variance of the weighted mean for the velocity error for non-variable sources (the expected behavior for non-variable sources with errors that are Normally distributed) and will have larger errors for stars that have evidence of variability. This combination method can be more robust to unidentified binary stars as stars with variation will have larger errors (Buttry et al., 2022). In principle, the center-of-mass velocity for binary stars should be used but this is extremely difficult to correctly identify with only a few velocity measurements. For repeat metallicity measurements we combine measurements with the weighted mean and variance of the weighted mean for the combined error. We have verified that using the best measurement (highest signal-to-noise) instead of the combined value gives similar results (when excluding binary stars).
@@ 2.2.1 Observations and Radial Velocity
The radial velocity was determined following Ji et al. (2020b). Each MIKE spectral order from 3700Å-6800Å was cross-correlated against a rest frame spectrum of the metal-poor red giant HD122563. As velocity uncertainties are dominated by wavelength calibration issues, we adopt the order-to-order scatter as the velocity uncertainty. The final velocity is km s-1. Separately, our IMACS data found PicII-1 to have velocities 330.0, 328.2, 329.9 km s-1, each with uncertainty 1 km s-1. There is no evidence for large radial velocity variations for PicII-1.
@@ 2.2.2 Stellar Parameters and Abundances
Stellar parameters were determined following the procedure in Frebel et al. (2013), i.e. excitation, ionization, and line strength balance of Fe lines with an empirical correction to match a photometric temperature scale. We find K, , , and for the effective temperature, surface gravity, microturbulence velocity, and metallicity, respectively. The uncertainties include statistical error from the Fe lines as well as systematic uncertainties of 150 K, 0.3 dex, and 0.2 for , , and respectively that reflect scatter in the Frebel et al. (2013) calibration. Line lists were adopted from Ji et al. (2020b). Abundances of most elements were determined from equivalent widths, but for the following features we determined the abundance with spectral synthesis: molecular CH bands; hyperfine structure in Sc, Mn, Co; and Al and Si lines somewhat blended with CH. We only use lines with Å because the S/N is too low at bluer wavelengths ( per pixel). The exception is a strong Al line at 3944Å as no other Al lines are available: this line is clearly detected, but given the S/N the abundance is extremely uncertain (constraint no better than 0.5 dex). We also determined formal upper limits for Sr, Ba, and Eu using spectral synthesis of the 4215Å, 4934Å, and 4129Å lines, respectively (see Ji et al. 2020a for more details).
Final abundances and uncertainties of PicII-1 are given in Table 2.2.2. For each element, is the number of lines measured. is the average abundance of those lines weighted by the abundance uncertainty. If and are the abundance and uncertainty of line , then we define and . The uncertainty for an individual line is found by propagating spectrum noise uncertainty. is the standard deviation of those lines (undefined if one line). accounts for the abundance uncertainty due to propagating individual line uncertainties, i.e. (McWilliam et al., 1995). [X/H] is the abundance relative to solar abundances from Asplund et al. (2009). [X/Fe] is calculated using either [Fe I/H] or [Fe II/H], depending on whether X is neutral or ionized. is the quadrature sum of , , and abundance uncertainties due to stellar parameter variations. is similar to , but when calculating the stellar parameter uncertainties we include variations in Fe. We use the difference in Fe I abundance for neutral species and the difference in Fe II abundance for ionized species to calculate this error. The [X/Fe] error is usually smaller than the [X/H] error, since abundance differences from changing and usually affect X and Fe in the same direction.
@@ 3.2 Spectroscopic Membership Classification
We assess the membership of each star based off its kinematics, chemistry, and photometric properties. We note that “background” stars here can refer to either foreground MW stars or LMC stars at a similar distance. The properties of the spectroscopic sample are summarized in Figure 1. We highlight our final spectroscopic sample, along with clear background stars, and background stars with a velocity similar to Pic II but classified as non-members. The velocity distribution of our spectroscopic sample has two peaks, a wide peak at which we associate with the MW and a second narrow peak at which we associate as Pic II. The majority of the stars closest to Pic II match the peak and the velocity association is further confirmed by selecting on proper motion. The proper motion selection can be done with the membership from Gaia proper motion based mixture models (Battaglia et al., 2022; Pace et al., 2022) or a proper motion and parallax selection such as the membership score from Tolstoy et al. (2023). All proper motion selections uncover the same velocity peak which confirms the Pic II systemic velocity.
@@
Of the non-members within of the Pic II velocity (green points in Figure 1), five stars are clear astrometric non-members and two stars have a metallicity (either CaT EW or CaHK colors) that is much larger than Pic II. Both of the metallicity outlier stars are consistent with the velocity and proper motion of Pic II and have . The first, (ID=10286300169398), has and is 1.2 dex more metal-rich than any of the Pic II members. This star does not have a in the MAGIC catalog but is located in the CaHK color space near this metallicity. Lastly, we note this star is located on the old, metal-poor stellar isochrone. While a [Fe/H] star in a UFD is not uncommon or implausible a large metallicity gap seems unlikely and we consider this star a likely LMC interloper. The second, (ID=10286300169327), does not have a CaT EW measurement and has . This star is redder than the stellar isochrone. As this star is even more metal-rich it is even less likely to be a Pic II member. In total, we identify 13 Pic II members, which includes two horizontal branch stars.
@@ 3.4 Stellar Kinematics
To assess the significance of resolving the velocity dispersion, we compute the Bayes factor () comparing a free velocity dispersion model and zero velocity dispersion model. The Bayes factor is the ratio of Bayesian evidence between two models and is a commonly utilized model selection tool (see Trotta, 2008, for a review). We use the pocoMC package (Karamanis et al., 2022a, b) to sample the posterior distributions and compute the Bayesian evidence for both models. We find strongly favoring the non-zero dispersion model101010The Bayes factor drops to 11.9 if we compare to the stellar only velocity dispersion of which still strongly favors the velocity dispersion model. For the linear prior we find comparing to the zero dispersion model.. We use Jeffreys’ scale to interpret the Bayes factor where values of , 1-2.5, 2.5-5, correspond to inconclusive, weak, moderate, and strong evidence in favor of one model, respectively (negative value favor the other model) (Trotta, 2008). Overall, we find strong evidence for a non-zero velocity dispersion.
@@ 4.1 The Nature of Pic II
Another defining characteristic of UFDs is their low neutron-capture element abundances from their inefficient star formation131313Although there are some exceptions: Reticulum II, (Ji et al., 2016a; Roederer et al., 2016); Tucana III, (Hansen et al., 2017; Marshall et al., 2019)), and Grus II (Hansen et al., 2020). (e.g., Koch et al., 2008; Frebel & Norris, 2015; Ji et al., 2019, 2020b). The neutron capture abundances for the brightest star in Pic II are consistent with this picture (Section 4.3). We only obtain upper limits for Ba, Eu, and Sr; the Sr and Ba upper limits exclude MW halo neutron capture values. Triangulum II and Grus I are examples of galaxies with unresolved or barely resolved velocity/metallicity dispersions, but whose low neutron-capture element content suggests they are (or once were) UFDs (Kirby et al., 2017; Venn et al., 2017; Ji et al., 2019). In contrast, all known globular clusters have moderately super-solar neutron-capture element abundances (e.g., Pritzl et al., 2005).
@@
Figure 5 shows the absolute V-band magnitude versus the metallicity and metallicity dispersion. Pic II is among the most metal-poor UFD known and is roughly 0.12 dex more metal-poor than Eridanus IV (Heiger et al., 2024), the previous most metal-poor UFD from spectroscopic measurements141414We note that Canes Venatici II and Hydra II have photometric CaHK metallicities that are more metal-poor or around the same metallicity as Pic II (Fu et al., 2023).. Interestingly, the LMC UFDs may be more metal-poor with smaller metallicity dispersions compared to MW UFDs with the caveats that there are a smaller number of LMC UFDs and the metallicity measurements come from an inhomogeneous sample. This may be due to lower star formation efficiencies of the LMC UFDs relative to the MW UFDs driven by their different environments. Our metallicity distribution measurements do not resolve a metallicity dispersion for Pic II.
We note that most dwarf galaxies contain RR Lyrae stars (e.g., Martínez-Vázquez, 2023; Tau et al., 2024) and we present an unsuccessful RRL search in Appendix A. We list the literature references for Figures 3-5 in Appendix B.
@@ 4.2 Orbit and Connection to the Magellanic Clouds
With a distance of kpc from the LMC, Drlica-Wagner et al. (2016) suggested that Pic II was likely accreted with the LMC and gravitationally bound to the LMC. With the 3D motions measured in this work, we compute the relative 3D velocity between the LMC and Pic II to be and an association is possible given the low relative velocity. Jethwa et al. (2016) built dynamical models to predict the phase space distribution of LMC and SMC UFD satellites. These models predict that Pic II will have a Galactic standard of rest line-of-sight velocity of and Pic II measured value is in the middle of this distribution.
@@
Pic II is the seventh LMC UFD to be confirmed with full phase space measurements and orbit modeling. The spatial distribution of the LMC UFDs is asymmetric and there are more LMC UFDs in the leading orbit of the LMC (Car II, Car III, Hyi I, Pic II, and Ret II) than trailing the orbit (Hor I and Phe II). This spatial anisotropy is similar to M31 (e.g., McConnachie & Irwin, 2006; Savino et al., 2022), where there are more satellites on the MW side of the satellite system compared to the far side. With the LMC system there are fewer confirmed UFDs overall and the system is being tidally heated by the MW. The satellite sensitivity is lower for more distant systems and the coverage for satellite searches is lower on the southern side of the LMC and this asymmetry may be due to selection effects. Of the known UFD candidates without spectroscopy, DELVE 2 (Cerny et al., 2021a) is a candidate member of the LMC system based on its spatial location and proper motion (Cerny et al., 2021a; Correa Magnus & Vasiliev, 2022; Vasiliev, 2024) and it is located on the far side of the LMC.

### v2

@@ Abstract
The advent of wide and deep optical surveys has precipitated the discovery of a large number of ultra-faint dwarf galaxies (UFDs, ) around the Milky Way (MW) (e.g., Willman et al., 2005; Belokurov et al., 2007; Laevens et al., 2015; Bechtol et al., 2015; Koposov et al., 2015a; Drlica-Wagner et al., 2015; Cerny et al., 2023a; Smith et al., 2023; Homma et al., 2024; Tan et al., 2025). UFDs are among the oldest, most chemically pristine, and dark matter dominated galaxies known (McConnachie, 2012; Simon, 2019; Pace, 2024) and are excellent probes to test CDM at small scales (Bullock & Boylan-Kolchin, 2017). UFDs are old and metal-poor, and their stars trace chemical enrichment from the first Population III stars (Simon & Geha, 2007; Brown et al., 2014; Frebel & Norris, 2015). The population of UFDs probes galaxy formation before the epoch of reionization (e.g., Bose et al., 2018) and constrains the small-scale power spectrum of dark matter (e.g., Jethwa et al., 2018; Newton et al., 2018; Nadler et al., 2021).
@@
The Magellanic SatelLites Survey111MagLiteS was a precursor survey to the DECam Local Volume Exploration survey. (MagLiteS; PI: K. Bechtol) was designed to search for additional satellites near the Magellanic Clouds (Drlica-Wagner et al., 2016), particularly on the southern side of the LMC/SMC that the DES did not cover. Three UFD candidates were discovered with MagLiteS: Pictor II (Pic II), Carina II (Car II), and Carina III (Car III) (Drlica-Wagner et al., 2016; Torrealba et al., 2016a). The radial velocities of Car II and Car III (Li et al., 2018a) suggested they were associated with the LMC based on simulated LMC satellites (Jethwa et al., 2016) and this association was confirmed with full 6-D phase space information from Gaia DR2 proper motion measurements (Gaia Collaboration et al., 2016, 2018) and orbit models that included the LMC (Erkal & Belokurov, 2020; Patel et al., 2020). Several other UFDs, Horologium I (Hor I), Hydrus I (Hyi I), Phoenix II (Phe II), and Reticulum II (Ret II), have additionally been associated with the LMC (Koposov et al., 2018; Simon, 2018; Kallivayalil et al., 2018; Erkal & Belokurov, 2020; Patel et al., 2020; Correa Magnus & Vasiliev, 2022; Battaglia et al., 2022; Pace et al., 2022; Vasiliev, 2023). While some of these UFDs may be SMC satellites, as the SMC is a satellite of the LMC, we only consider the LMC connection in this work.
The LMC UFD population enables studying UFDs that formed in a different and lower mass environment compared to the in-situ MW UFDs. Star formation history measurements suggest that on average the LMC satellites were quenched more recently compared to MW satellites (Sacchi et al., 2021; Durbin et al., 2025). This may be due to patchy reionization allowing lower density environments to form stars longer (e.g., Lunnan et al., 2012; Katz et al., 2020; Kim et al., 2023). Conversely, no difference has been seen in the size-luminosity plane between the LMC and MW UFDs (Richstein et al., 2024) but the number of LMC confirmed UFDs is an order of magnitude smaller than MW UFDs and confirming more LMC UFDs is key to studying environmental differences.
@@ 2.1.2 Velocity and Metallicity Measurements
For our primary analysis, we combined repeat measurements at the catalog level. For the combined velocities we fit a free velocity and dispersion to the repeat measurements and use the mean velocity and error for the star’s intrinsic properties. The velocity errors from this method will match the variance of the weighted mean for non-variable stars (the expected behavior for non-variable sources with errors that are Normally distributed) and will be larger for stars with evidence of variability. This combination method can be more robust to unidentified binary stars as stars with variation will have larger errors (Buttry et al., 2022). In principle, the center-of-mass velocity for binary stars should be used but this is extremely difficult to correctly identify with only a few velocity measurements. For repeat metallicity measurements we combine measurements with the weighted mean and variance of the weighted mean for the combined error. We have verified that using the best measurement (highest signal-to-noise) instead of the combined value gives similar results (when excluding binary stars).
@@ 2.2.1 Observations and Radial Velocity
The radial velocity was determined following Ji et al. (2020a). Each MIKE spectral order from 3700Å-6800Å was cross-correlated against a rest frame spectrum of the metal-poor red giant HD122563. As velocity uncertainties are dominated by wavelength calibration issues, we adopt the order-to-order scatter as the velocity uncertainty. The final velocity is km s-1. Separately, our IMACS data found PicII-1 to have velocities 330.0, 328.2, 329.9 km s-1, each with uncertainty 1 km s-1. There is no evidence for large radial velocity variations for PicII-1.
@@ 2.2.2 Stellar Parameters and Abundances
Stellar parameters were determined following the procedure in Frebel et al. (2013), i.e. excitation, ionization, and line strength balance of Fe lines with an empirical correction to match a photometric temperature scale. We find K, , , and for the effective temperature, surface gravity, microturbulence velocity, and metallicity, respectively. The uncertainties include statistical error from the Fe lines as well as systematic uncertainties of 150 K, 0.3 dex, and 0.2 for , , and respectively that reflect scatter in the Frebel et al. (2013) calibration. Line lists were adopted from Ji et al. (2020a). Abundances of most elements were determined from equivalent widths, but for the following features we determined the abundance with spectral synthesis: molecular CH bands; hyperfine structure in Sc, Mn, Co; and Al and Si lines somewhat blended with CH. We only use lines with Å because the S/N is too low at bluer wavelengths ( per pixel). The exception is a strong Al line at 3944Å as no other Al lines are available: this line is clearly detected, but given the S/N the abundance is extremely uncertain (constraint no better than 0.5 dex). We also determined formal upper limits for Sr, Ba, and Eu using spectral synthesis of the 4215Å, 4934Å, and 4129Å lines, respectively (see Ji et al. 2020b for more details).
Final abundances and uncertainties of PicII-1 are given in Table 2.2.2. For each element, is the number of lines measured. is the average abundance of those lines weighted by the abundance uncertainty. If and are the abundance and uncertainty of line , then we define and . The uncertainty for an individual line is found by propagating spectrum noise uncertainty. is the standard deviation of those lines (undefined if one line). accounts for the abundance uncertainty due to propagating individual line uncertainties, i.e. (McWilliam et al., 1995). [X/H] is the abundance relative to solar abundances from Asplund et al. (2009). [X/Fe] is calculated using either [Fe I/H] or [Fe II/H], depending on whether X is neutral or ionized. is the quadrature sum of , , and abundance uncertainties due to stellar parameter variations. is similar to , but when calculating the stellar parameter uncertainties we include variations in Fe. We use the difference in Fe I abundance for neutral species and the difference in Fe II abundance for ionized species to calculate this error. The [X/Fe] error is usually smaller than the [X/H] error, since abundance differences from changing and usually affect X and Fe in the same direction. All analyses are conducted in LTE, but non-LTE effects are significant for many elements (e.g., Sitnova et al., 2021). Due to inhomogeneities in how non-LTE corrections are determined and applied, we reserve a full discussion of this for a future analysis of all UFD stars. However, the most significant effects in cool giants are that Al and Mn increase by 0.5-1.0 dex (Nordlander & Lind, 2017; Bergemann et al., 2019); Na and K decrease by up to 0.5 dex (Lind et al., 2011; Reggiani et al., 2019). Other elements change by dex (Sitnova et al., 2021).
@@ 3.2 Spectroscopic Membership Classification
We assess the membership of each star based off its kinematics, chemistry, and photometric properties. We note that “background” stars here can refer to either foreground MW stars or LMC stars at a similar distance. The properties of the spectroscopic sample are summarized in Figure 1. We highlight our final spectroscopic sample, along with clear background stars, and background stars with a velocity similar to Pic II but classified as non-members. The velocity distribution of our spectroscopic sample has two peaks, a wide peak at which we associate with the MW and a second narrow peak at which we associate as Pic II. The majority of the stars closest to Pic II match the peak and the velocity association is further confirmed by selecting on proper motion. The proper motion selection can be done with the membership from Gaia proper motion based mixture models (membership probability ; Battaglia et al., 2022; Pace et al., 2022) or a proper motion and parallax selection such as the membership score () from Tolstoy et al. (2023). All proper motion selections uncover the same velocity peak which confirms the Pic II systemic velocity.
@@
Of the non-members within of the Pic II velocity (green points in Figure 1), five stars are clear astrometric non-members and two stars have a metallicity (either CaT EW or CaHK colors) that is much larger than Pic II. Both of the metallicity outlier stars are consistent with the velocity and proper motion of Pic II and have . The first, (ID=10286300169398), has and is 1.2 dex more metal-rich than any of the Pic II members. This star does not have a in the MAGIC catalog but is located in the CaHK color space near this metallicity. Lastly, we note this star is located on the old, metal-poor stellar isochrone. While a [Fe/H] star in a UFD is not uncommon or implausible a large metallicity gap seems unlikely and we consider this star a likely LMC interloper. The second, (ID=10286300169327), does not have a CaT EW measurement and has . This star is redder than the stellar isochrone. As this star is even more metal-rich it is even less likely to be a Pic II member. While both these stars are more metal-poor than the bulk of LMC, they are consistent with the LMC halo metallicity distribution (e.g., Borissova et al., 2006). In total, we identify 13 Pic II members, which includes two horizontal branch stars.
@@ 3.4 Stellar Kinematics
To assess the significance of resolving the velocity dispersion, we compute the Bayes factor () comparing a free velocity dispersion model and zero velocity dispersion model. The Bayes factor is the ratio of Bayesian evidence between two models and is a commonly utilized model selection tool (see Trotta, 2008, for a review). We use the pocoMC package (Karamanis et al., 2022b, a) to sample the posterior distributions and compute the Bayesian evidence for both models. We find strongly favoring the non-zero dispersion model101010The Bayes factor drops to 11.9 if we compare to the stellar only velocity dispersion of which still strongly favors the velocity dispersion model. For the linear prior we find comparing to the zero dispersion model.. We use Jeffreys’ scale to interpret the Bayes factor where values of , 1-2.5, 2.5-5, correspond to inconclusive, weak, moderate, and strong evidence in favor of one model, respectively (negative value favor the other model) (Trotta, 2008). Overall, we find strong evidence for a non-zero velocity dispersion.
@@ 4.1 The Nature of Pic II
Another defining characteristic of UFDs is their low neutron-capture element abundances from their inefficient star formation131313Although there are some exceptions: Reticulum II, (Ji et al., 2016a; Roederer et al., 2016); Tucana III, (Hansen et al., 2017; Marshall et al., 2019)), and Grus II (Hansen et al., 2020). (e.g., Koch et al., 2008; Frebel & Norris, 2015; Ji et al., 2019, 2020a). The neutron capture abundances for the brightest star in Pic II are consistent with this picture (Section 4.3). We only obtain upper limits for Ba, Eu, and Sr; the Sr and Ba upper limits exclude MW halo neutron capture values. Triangulum II and Grus I are examples of galaxies with unresolved or barely resolved velocity/metallicity dispersions, but whose low neutron-capture element content suggests they are (or once were) UFDs (Kirby et al., 2017; Venn et al., 2017; Ji et al., 2019). In contrast, all known globular clusters have moderately super-solar neutron-capture element abundances (e.g., Pritzl et al., 2005).
@@
Figure 5 shows the absolute V-band magnitude versus the metallicity and metallicity dispersion. Pic II is among the most metal-poor UFD known and is roughly 0.12 dex more metal-poor than Eridanus IV (Heiger et al., 2024), the previous most metal-poor UFD from spectroscopic measurements141414We note that Canes Venatici II and Hydra II have photometric CaHK metallicities that are more metal-poor or around the same metallicity as Pic II (Fu et al., 2023) but these photometric metallicity measurements are more metal-poor than previous spectroscopic measurements (Kirby et al., 2013b, 2015).. The LMC UFDs are more metal-poor than expected from the stellar mass-metallicity relationship and Pic II is 0.44 dex more metal-poor than expected from this relation. Our metallicity measurements do not resolve a metallicity dispersion for Pic II and our upper limit is smaller than most UFD measurements. Only Hor I, Eridanus IV, Leo VI, and Tucana III have metallicity dispersions around or smaller than 0.20 dex, similar to the Pic II upper limit (Koposov et al., 2015b; Simon et al., 2017; Heiger et al., 2024; Tan et al., 2025). The small metallicity dispersion of Pic II may be due to its lower star formation efficiency relative to other UFDs.
We note that most dwarf galaxies contain RR Lyrae stars (e.g., Mart´ınez-Vázquez, 2023; Tau et al., 2024) and we present an unsuccessful RRL search in Appendix A. We list the literature references for Figures 3-5 in Appendix B.
@@ 4.2 Orbit and Connection to the Magellanic Clouds
With a 3D distance of kpc from the LMC, Drlica-Wagner et al. (2016) suggested that Pic II was likely accreted with the LMC and gravitationally bound to the LMC. With the 3D motions measured in this work, we compute the relative 3D velocity between the LMC and Pic II to be and an association is possible given the low relative velocity. Jethwa et al. (2016) built dynamical models to predict the phase space distribution of LMC and SMC UFD satellites. These models predict that Pic II will have a Galactic standard of rest line-of-sight velocity of and Pic II measured value is in the middle of this distribution.
@@
Pic II is the seventh LMC UFD to be confirmed with full phase space measurements and orbit modeling. The spatial distribution of the LMC UFDs is asymmetric and there are more LMC UFDs in the leading orbit of the LMC (Car II, Car III, Hyi I, Pic II, and Ret II) than trailing the orbit (Hor I and Phe II). This spatial anisotropy is similar to M31 (e.g., McConnachie & Irwin, 2006; Savino et al., 2022), where there are more satellites on the MW side of the satellite system compared to the far side. With the LMC system there are fewer confirmed UFDs overall and the system is being tidally heated by the MW. The satellite sensitivity is lower for more distant systems and the coverage for satellite searches is lower on the southern side of the LMC and this asymmetry may be due to selection effects. This asymmetry does not appear long lived as the number of LMC satellites on the near/far sides changes multiple times during our orbital integration. Of the known UFD candidates without spectroscopy, DELVE 2 (Cerny et al., 2021a) is a candidate member of the LMC system based on its spatial location and proper motion (Cerny et al., 2021a; Correa Magnus & Vasiliev, 2022; Vasiliev, 2024) and it is located on the far side of the LMC.